# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 2.4 Unpolarized background
$\delta\mu^{*}_{A}$ $=\dfrac{\left[\frac{2\omega_{*}}{T^{2}}(A\sigma^{\omega}_{A}-B\sigma^{\omega}_{H})+\frac{i}{3\tau_{H}}A\right]\kappa_{\Omega}\delta\mu^{*}_{V}}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}},$
$\delta\mu^{*}_{H}$ $=\dfrac{\left[\frac{2\omega_{*}}{T^{2}}(B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H})+\frac{i}{3\tau_{A}}B\right]\kappa_{\Omega}\delta\mu^{*}_{V}}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}}.$ (38)
The previous solutions are valid so long as the denominator does not vanish for the particular mode $\omega_{*}$ Looking ahead, we note that we will encounter the opposite situation in Subsec. 3.2). Then, it is more convenient to solve $\mathbb{M}_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}=0$ for the first and second rows ($\ell=V,A$) and take $\delta\mu_{H}$ as the reference amplitude. The relations between the amplitudes are given by:
@@
We stress that Eqs. (38) and (39) are equivalent for a given mode $\omega_{*}$ provided the denominators do not vanish.
Before ending this section, we quote from Ref. HVWp1 the solutions for the angular velocities satisfying Eq. (37) in the case of conserved axial and helical charges, corresponding to the $\tau_{A},\tau_{H}\rightarrow\infty$ limit:
$\omega_{\rm im.}=0,\qquad\omega_{h}^{\pm}=\pm\frac{\kappa_{\Omega}T^{2}}{2}\sqrt{\frac{\sigma^{\omega}_{A}(A^{2}+B^{2})-2AB\sigma^{\omega}_{H}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}}.$ (40)
@@
$\delta\mu^{h;\pm}_{A}$ $=\dfrac{T^{2}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\kappa_{\Omega}}{2\omega_{h}^{\pm}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\delta\mu^{h;\pm}_{V}\,,$ $\delta\mu_{H}$ $=\dfrac{T^{2}(B\sigma_{A}^{\omega}-A\sigma_{H}^{\omega})\kappa_{\Omega}}{2\omega_{h}^{\pm}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\delta\mu^{h;\pm}_{V}\,.$ (42)
In Sec. 3, we shall study Eqs. (37) and (38) for the cases when $0<\tau_{H}<\infty$ (the helical charge has a finite relaxation time and thus it is not conserved) and $\tau_{A}\rightarrow\infty$ (the axial charge is conserved). The case with finite $\tau_{A}$ will be studied in Sec. 4. In both cases, we will take various limits, such as the high temperature or the degenerate gas limit, discussing the effect of charge non-conservation on the wave spectrum uncovered in the case of the VAH plasma with conserved charges, discussed in Ref. HVWp1 .
@@ 3 Non-conserved helical charge
As in the previous section, we will consider a plasma with vanishing helical and axial background (mean) chemical potentials. The assumption of vanishing helical chemical potential is justified by the fact that at finite values of $\tau_{H}$, the violation of the conservation of helical charge prevents introducing the helical chemical as a Lagrange multiplier in the grand canonical ensemble. The assumption of the vanishing $\mu_{A}$ background is not imposed physically but is technically convenient for the simplified analysis. The generalization of our discussion in this section to the case of finite $\mu_{A}$ is relatively straightforward. All in all, Eq. (9) becomes
$\nabla_{\mu}J^{\mu}_{V}=0\,,\qquad\nabla_{\mu}J^{\mu}_{A}=0\,,\qquad\nabla_{\mu}J^{\mu}_{H}=-\frac{T^{2}}{3\tau_{H}}\delta\mu_{H}\,.$ (43)
Physically, the non-conservation of helicity in QED arises from helicity-violating pair-annihilation processes of the form $e_{R}^{+}e_{L}^{-}\to e_{L}^{+}e_{R}^{-}$. Each such process violates helicity by two units while preserving chirality. Analogous processes are mediated by gluon interchange in QCD. For small helical imbalance, the helical relaxation time arising from such processes was estimated for QCD in the weakly-interacting regime as Ambrus:2019khr  [^6] The numerical value quoted here differs from the one reported in Ref. Ambrus:2019khr due to a factor of two error when going from Eq. (160) to Eq. (167) – see Erratum Helicity:2024err for details.
$\tau_{H}^{\rm QCD}\simeq\left(\dfrac{250\ {\rm MeV}}{k_{B}T}\right)\left(\dfrac{1}{\alpha_{\rm QCD}}\right)\left(\dfrac{2}{N_{f}}\right)\times 4.8\ {\rm fm}/c.$ (44)
@@
In the case $\tau_{A}\rightarrow\infty$, the matrix $\mathbb{M}/T^{2}$ in Eq. (34) reduces to
@@
$\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3\tau_{H}}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)-\frac{\kappa_{\Omega}^{2}A^{2}}{H}\right]=0.$ (46)
@@
$\delta\mu^{*}_{A}$ $=\dfrac{\left(A\sigma^{\omega}_{A}-B\sigma^{\omega}_{H}+\frac{iT^{2}A}{6\omega_{*}\tau_{H}}\right)\kappa_{\Omega}\delta\mu^{*}_{V}}{\frac{2\omega_{*}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{i}{3\tau_{H}}\sigma^{\omega}_{A}},$ $\delta\mu^{*}_{H}$ $=\dfrac{\left(B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H}\right)\kappa_{\Omega}\delta\mu_{V}^{*}\,}{\frac{2\omega_{*}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{i}{3\tau_{H}}\sigma^{\omega}_{A}}\,.$ (47)
@@ 3.1 Small vector chemical potential limit
$\omega_{h;0}^{\pm}=-\frac{i}{2\tau_{H}}\pm kc_{h}\sqrt{1-\frac{k_{\rm th}^{2}}{k^{2}}},\qquad k_{\rm th}=\frac{1}{2\tau_{H}c_{h}}=\frac{\pi^{2}}{12\ln 2}\frac{T}{\Omega\tau_{H}}.$ (52)
Notice that we work with a dissipative (lossy) medium so that the apparent (in fact, infinite) superluminal group velocity $v_{\rm gr.}=\partial{\rm Re}\,[\omega(k)]/\partial k$ at the merging point $k=k_{\rm th}$ does not violate causality as long as the asymptotic causality condition at large momentum, $k\to\infty$, is fulfilled Pu:2009fj .
@@
At first order with respect to $\alpha_{V}$, Eq. (46) reads
@@
$-\frac{4}{3}\left(\frac{\kappa_{\Omega}\ln 2}{\pi^{2}}\right)^{2}\left[\omega_{{\mathfrak{a}};2}+\frac{i}{49\tau_{H}(\ln 2)^{2}}\right]=0\quad\Rightarrow\quad\omega_{{\mathfrak{a}};2}=-\frac{i}{49\tau_{H}(\ln 2)^{2}}.$ (54)
@@
$\delta\mu_{V}^{\mathfrak{a}}\simeq-\frac{i}{k\tau_{H}c_{h}}\frac{\alpha_{V}}{7\ln 2}\delta\mu_{A}^{\mathfrak{a}},\qquad\delta\mu^{\mathfrak{a}}_{H}\simeq-\frac{\alpha_{V}}{7\ln 2}\delta\mu^{\mathfrak{a}}_{A}.$ (55)
@@
$\delta\mu^{h;\pm}_{H}$ $\simeq\frac{\omega^{\pm}_{h;0}}{kc_{h}}\delta\mu_{V}^{h;\pm}=\left(-\frac{ik_{\rm th}}{k}\pm c_{h}\sqrt{k^{2}-k_{\rm th}^{2}}\right)\delta\mu^{h;\pm}_{V},$
@@
$=\left[\frac{i}{\tau_{H}}\left(\frac{\pi^{2}}{84(\ln 2)^{2}}+1\right)\pm 2c_{h}\sqrt{k^{2}-k_{\rm th}^{2}}\left(\frac{\pi^{2}}{84(\ln 2)^{2}}-1\right)\right]\frac{\alpha_{V}}{\kappa_{\Omega}}\delta\mu^{\pm}_{V}.$ (56)
@@
$\bar{\mu}_{V}(0,z)=\mu_{V}+\delta\mu_{V}\cos(kz),\qquad\bar{\mu}_{A}(0,z)=\bar{\mu}_{H}(0,z)=0.$ (57)
To leading order in $\alpha_{V}$, this setup assumes that the amplitudes of the axial mode vanish, $\delta\mu_{\ell}^{\mathfrak{a}}\simeq 0$. The amplitudes $\delta\mu_{V}^{h;\pm}$ and $\delta\mu_{H}^{h;\pm}$ corresponding to the vector-helical modes satisfy
$\delta\mu_{V}^{h;\pm}\simeq\frac{\delta\mu_{V}}{2}\left(1\pm\frac{ik_{\rm th}}{\sqrt{k^{2}-k_{\rm th}^{2}}}\right),\qquad\delta\mu_{H}^{h;\pm}\simeq\pm\frac{k^{2}\tau_{H}c_{h}\delta\mu_{V}}{2\sqrt{k^{2}-k_{\rm th}^{2}}},$ (58)
while $\delta\mu_{A}^{h;\pm}$ is proportional to $\alpha_{V}$, as shown in Eq. (56). The full solution propagates as
$\delta\bar{\mu}_{V}(t,z)$ $\simeq\delta\mu_{V}e^{-t/2\tau_{H}}\cos(kz)\left[\cos\left(c_{h}t\sqrt{k^{2}-k_{\rm th}^{2}}\right)+\frac{\sin\left(c_{h}t\sqrt{k^{2}-k^{2}_{\rm th}}\right)}{2\tau_{H}c_{h}\sqrt{k^{2}-k_{\rm th}^{2}}}\right],$
$\delta\bar{\mu}_{H}(t,z)$ $\simeq\frac{k^{2}\tau_{H}c_{h}\delta\mu_{V}}{\sqrt{k^{2}-k_{\rm th}^{2}}}e^{-t/2\tau_{H}}\cos(kz)\sin\left(c_{h}t\sqrt{k^{2}-k_{\rm th}^{2}}\right).$ (59)
@@
$\delta\bar{\mu}_{V}{\biggl{|}}_{k<k_{\rm th}}$ $\rightarrow\frac{\delta\mu_{V}}{2}\left(1+\frac{k_{\rm th}}{\sqrt{k_{\rm th}^{2}-k^{2}}}\right)e^{-\frac{t}{2\tau_{H}}(1-\sqrt{1-k^{2}/k_{\rm th}^{2}})}\cos(kz),$
$\delta\bar{\mu}_{H}{\biggl{|}}_{k<k_{\rm th}}$ $\rightarrow\frac{\delta\mu_{V}k^{2}\tau_{H}c_{h}}{2\sqrt{k_{\rm th}^{2}-k^{2}}}e^{-\frac{t}{2\tau_{H}}(1-\sqrt{1-k^{2}/k_{\rm th}^{2}})}\cos(kz).$ (60)
@@
Refer to caption Refer to caption
Figure 1: Effect of the helicity relaxation time $\tau_{H}$ on the propagation of the initial Gaussian perturbation of the vector charge, shown in Eq. (61), in the case of conserved axial charge ($\tau_{A}\rightarrow\infty$). The profiles correspond to various values of $\tau_{H}$ and are represented with respect to $z/\sigma$, where $\sigma$ is the Gaussian width. The background fluid is neutral and unpolarized ($\mu_{V}=\mu_{A}=\mu_{H}=0$), and the temperature and angular velocity are set for definiteness to $T=300$ MeV and $\Omega=6.6$ MeV, respectively.
@@
In the limit when the helicity charge is conserved, $\tau_{H}\rightarrow\infty$, the square brackets on the second line of Eq. (62) evaluate to $\cos(c_{h}kt)$ and the integral with respect to $k$ can be performed analytically (see also Eq. (3.36) of Ref. HVWp1 ):
$\lim_{\tau_{H}\rightarrow\infty}\delta\bar{\mu}_{V}(t,z)=\frac{\delta\bar{\mu}_{V;0}}{2}\left[e^{-(z-c_{h}t)^{2}/2\sigma^{2}}+e^{-(z+c_{h}t)^{2}/2\sigma^{2}}\right],$
$\lim_{\tau_{H}\rightarrow\infty}\delta\bar{\mu}_{H}(t,z)=\frac{\delta\bar{\mu}_{V;0}}{2}\left[e^{-(z-c_{h}t)^{2}/2\sigma^{2}}-e^{-(z+c_{h}t)^{2}/2\sigma^{2}}\right].$ (63)
In Fig. 1, we represented $\delta\bar{\mu}_{V}(t,z)$, $\delta\bar{\mu}_{H}(t,z)$ and $\delta\bar{\mu}_{A}(t,z)$ with respect to $z/\sigma$, when $c_{h}t/\sigma=5$, for the case of a neutral background ($\mu_{V}=\mu_{A}=\mu_{H}=0$) and temperature $T=300$ MeV. We considered various values of $c_{h}\tau_{H}/\sigma$. It can be seen that, compared to the case when the helicity charge is conserved and $\tau_{H}\rightarrow\infty$ (shown with the black line), the propagation features are damped as $\tau_{H}$ is decreased. At the lowest value of $c_{h}\tau_{H}/\sigma=0.5$, $\delta\mu_{V}(t,z)$ simply decays on the time scale given by $\tau_{H}$.

### v2

@@ 2.4 Unpolarized background
$\delta\mu^{*}_{A}=$
$\dfrac{\left[\frac{2\omega_{*}}{T^{2}}(A\sigma^{\omega}_{A}-B\sigma^{\omega}_{H})+\frac{i}{3\tau_{H}}A\right]\kappa_{\Omega}\delta\mu^{*}_{V}}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}},$
$\delta\mu^{*}_{H}=$
$\dfrac{\left[\frac{2\omega_{*}}{T^{2}}(B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H})+\frac{i}{3\tau_{A}}B\right]\kappa_{\Omega}\delta\mu^{*}_{V}}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}}.$ (34)
The previous solutions are valid so long as the denominator does not vanish for the particular mode $\omega_{*}$. Looking ahead, in Subsec. 3.2 we will encounter the situation when the denominator does vanish. Then, it is more convenient to solve $\mathbb{M}_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}=0$ for the first and second rows ($\ell=V,A$) and take $\delta\mu_{H}$ as the reference amplitude. The relations between the amplitudes are given by:
@@
Note that the above relations contain an implicit dependence on the helicity relaxation time, $\tau_{H}$, through the mode angular frequency, $\omega_{*}$. We stress that Eqs. (34) and (35) are equivalent for a given mode $\omega_{*}$ provided the denominators do not vanish.
Before ending this section, we quote from Ref. Morales-Tejera:2024uzg the solutions for the angular velocities satisfying Eq. (33) in the case of conserved axial and helical charges, corresponding to the $\tau_{A},\tau_{H}\rightarrow\infty$ limit:
$\omega_{\rm im.}=0,$
$\omega_{h}^{\pm}=\pm\frac{\kappa_{\Omega}T^{2}}{2}\sqrt{\frac{\sigma^{\omega}_{A}(A^{2}+B^{2})-2AB\sigma^{\omega}_{H}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}}.$ (36)
@@
$\delta\mu^{h;\pm}_{A}$ $=\dfrac{T^{2}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\kappa_{\Omega}}{2\omega_{h}^{\pm}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\delta\mu^{h;\pm}_{V}\,,$
$\delta\mu_{H}$ $=\dfrac{T^{2}(B\sigma_{A}^{\omega}-A\sigma_{H}^{\omega})\kappa_{\Omega}}{2\omega_{h}^{\pm}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\delta\mu^{h;\pm}_{V}\,.$ (38)
In Sec. 3, we shall study Eqs. (33) and (34) for the cases when $0<\tau_{H}<\infty$ (the helical charge has a finite relaxation time and thus it is not conserved) and $\tau_{A}\rightarrow\infty$ (the axial charge is conserved). The case with finite $\tau_{A}$ will be studied in Sec. 4. In both cases, we will take various limits, such as the high temperature or the degenerate gas limit, discussing the effect of charge non-conservation on the wave spectrum uncovered in the case of the plasma with conserved V/A/H charges, discussed in Ref. Morales-Tejera:2024uzg .
@@ 3 Non-conserved helical charge
As in the previous section, we will consider a plasma with vanishing helical and axial background (mean) chemical potentials. The assumption of vanishing helical chemical potential is justified by the fact that at finite values of $\tau_{H}$, the violation of the conservation of helical charge prevents introducing the helical chemical potential as a Lagrange multiplier in the grand canonical ensemble. The assumption of the vanishing $\mu_{A}$ background is not imposed physically but is technically convenient to simplify the analysis. The generalization of our discussion in this section to the case of finite $\mu_{A}$ is relatively straightforward. All in all, Eq. (9) becomes
$\partial_{\mu}J^{\mu}_{V}=0\,,\ \ \partial_{\mu}J^{\mu}_{A}=0\,,\ \ \partial_{\mu}J^{\mu}_{H}=-\frac{T^{2}}{3\tau_{H}}\delta\mu_{H}\,.$ (39)
Physically, the non-conservation of helicity in QED arises from helicity-violating pair-annihilation processes of the form $e_{R}^{+}e_{L}^{-}\to e_{L}^{+}e_{R}^{-}$. Each such process violates helicity by two units while preserving chirality. Analogous processes are mediated by gluon interchange in QCD. For small helical imbalance, the helical relaxation time arising from such processes was estimated for QCD in the weakly-interacting regime as Ambrus:2019khr ; Helicity:2024err
$\tau_{H}^{\rm QCD}\simeq\left(\dfrac{250\ {\rm MeV}}{k_{B}T}\right)\left(\dfrac{1}{\alpha_{\rm QCD}}\right)^{2}\left(\dfrac{2}{N_{f}}\right)\\ \times 4.8\ {\rm fm}/c.$ (40)
@@
In the case $\tau_{A}\rightarrow\infty$, the matrix $\mathbb{M}/T^{2}$ in Eq. (15a) reduces to
@@
$\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right.\\ \left.-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3\tau_{H}}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)-\frac{\kappa_{\Omega}^{2}A^{2}}{H}\right]=0.$ (42)
@@
$\delta\mu^{*}_{A}$ $=\dfrac{\left(A\sigma^{\omega}_{A}-B\sigma^{\omega}_{H}+\frac{iT^{2}A}{6\omega_{*}\tau_{H}}\right)\kappa_{\Omega}\delta\mu^{*}_{V}}{\frac{2\omega_{*}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{i}{3\tau_{H}}\sigma^{\omega}_{A}},$
$\delta\mu^{*}_{H}$ $=\dfrac{\left(B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H}\right)\kappa_{\Omega}\delta\mu_{V}^{*}\,}{\frac{2\omega_{*}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{i}{3\tau_{H}}\sigma^{\omega}_{A}}\,.$ (43)
@@ 3.1 Small vector chemical potential limit
$\omega_{h;0}^{\pm}=-\frac{i}{2\tau_{H}}\pm kc_{h}\sqrt{1-\frac{k_{\rm th}^{2}}{k^{2}}},$
$k_{\rm th}=\frac{1}{2\tau_{H}c_{h}}=\frac{\pi^{2}}{12\ln 2}\frac{T}{\Omega\tau_{H}}.$ (48)
The above relation provides the dispersion relation for the HVW in a neutral plasma, which is fully consistent with the previously reported result for the HVW in the same limit [c.f. Eq. (118) in Ref.Ambrus:2019khr ]. Notice that we work with a dissipative (lossy) medium so that the apparent (in fact, infinite) superluminal group velocity $v_{\rm gr.}=\partial{\rm Re}\,[\omega(k)]/\partial k$ at the merging point $k=k_{\rm th}$ does not violate causality as long as the asymptotic causality condition at large momentum, $k\to\infty$, is fulfilled Pu:2009fj .
@@
At first order with respect to $\alpha_{V}$, Eq. (42) gives
@@
$-\frac{4}{3}\left(\frac{\kappa_{\Omega}\ln 2}{\pi^{2}}\right)^{2}\left[\omega_{{\mathfrak{a}};2}+\frac{i}{\tau_{H}(7\ln 2)^{2}}\right]=0$
$\Rightarrow\quad\omega_{{\mathfrak{a}};2}=-\frac{i}{\tau_{H}(7\ln 2)^{2}}.$ (50)
@@
$\delta\mu_{V}^{\mathfrak{a}}$ $\simeq-\frac{i}{k\tau_{H}c_{h}}\frac{\alpha_{V}}{7\ln 2}\delta\mu_{A}^{\mathfrak{a}}\,,$ (51a)
$\delta\mu^{\mathfrak{a}}_{H}$ $\simeq-\frac{\alpha_{V}}{7\ln 2}\delta\mu^{\mathfrak{a}}_{A}\,.$ (51b)
@@
$\delta\mu^{h;\pm}_{H}$ $\simeq\frac{\omega^{\pm}_{h;0}}{kc_{h}}\delta\mu_{V}^{h;\pm}$
$=\left(-ik_{\rm th}\pm\sqrt{k^{2}-k_{\rm th}^{2}}\right)\frac{\delta\mu^{h;\pm}_{V}}{k},$
@@
$=\left[\frac{i}{\tau_{H}}\left(\frac{\pi^{2}}{84(\ln 2)^{2}}+1\right)\right.$
±$\left.\pm 2c_{h}\sqrt{k^{2}-k_{\rm th}^{2}}\left(\frac{\pi^{2}}{84(\ln 2)^{2}}-1\right)\right]\frac{\alpha_{V}}{\kappa_{\Omega}}\delta\mu^{\pm}_{V}.$ (52)
@@
$\bar{\mu}_{V}(0,z)=\mu_{V}+\delta\bar{\mu}_{V;0}\cos(kz),$
$\bar{\mu}_{A}(0,z)=\bar{\mu}_{H}(0,z)=0.$ (53)
To leading order in $\alpha_{V}$, this setup assumes that the amplitudes in the axial mode vanish, $\delta\mu_{\ell}^{\mathfrak{a}}\simeq 0$. The amplitudes $\delta\mu_{V}^{h;\pm}$ and $\delta\mu_{H}^{h;\pm}$ corresponding to the vector-helical modes satisfy
$\delta\mu_{V}^{h;\pm}$ $\simeq\frac{\delta\bar{\mu}_{V;0}}{2}\left(1\pm\frac{ik_{\rm th}}{\sqrt{k^{2}-k_{\rm th}^{2}}}\right),$
$\delta\mu_{H}^{h;\pm}$$\simeq\pm\frac{k\delta\bar{\mu}_{V;0}}{2\sqrt{k^{2}-k_{\rm th}^{2}}},$ (54)
while $\delta\bar{\mu}_{A;0}^{h;\pm}$ is proportional to $\alpha_{V}$, as shown in Eq. (52). The full solution reads:
$\delta\bar{\mu}_{V}(t,z)$ $\simeq\delta\bar{\mu}_{V;0}e^{-t/2\tau_{H}}\cos(kz)$
×$\times\left[\cos\left(c_{h}t\sqrt{k^{2}-k_{\rm th}^{2}}\right)+\frac{\sin\left(c_{h}t\sqrt{k^{2}-k^{2}_{\rm th}}\right)}{2\tau_{H}c_{h}\sqrt{k^{2}-k_{\rm th}^{2}}}\right],$
$\delta\bar{\mu}_{H}(t,z)$ $\simeq\frac{k\delta\bar{\mu}_{V;0}e^{-t/2\tau_{H}}}{\sqrt{k^{2}-k_{\rm th}^{2}}}$
×$\times\sin(kz)\sin\left(c_{h}t\sqrt{k^{2}-k_{\rm th}^{2}}\right).$ (55)
@@
$\delta\bar{\mu}_{V}{\biggl{\rvert}}_{k<k_{\rm th}}$ $\rightarrow\frac{\delta\bar{\mu}_{V;0}}{2}\left(1+\frac{k_{\rm th}}{\sqrt{k_{\rm th}^{2}-k^{2}}}\right)\times$
×$\times e^{-\frac{t}{2\tau_{H}}(1-\sqrt{1-k^{2}/k_{\rm th}^{2}})}\cos(kz),$
$\delta\bar{\mu}_{H}{\biggl{\rvert}}_{k<k_{\rm th}}$ $\rightarrow\frac{\delta\bar{\mu}_{V;0}k^{2}\tau_{H}c_{h}}{2\sqrt{k_{\rm th}^{2}-k^{2}}}$
×$\times e^{-\frac{t}{2\tau_{H}}(1-\sqrt{1-k^{2}/k_{\rm th}^{2}})}\sin(kz).$ (56)
@@
Figure 1: Effect of the helicity relaxation time $\tau_{H}$ on the propagation of the initial Gaussian perturbation of the vector charge, shown in Eq. (57), in the case of conserved axial charge ($\tau_{A}\rightarrow\infty$). The profiles correspond to various values of $\tau_{H}$ and are represented with respect to $z/\sigma$, where $\sigma$ is the Gaussian width. The background fluid is neutral and unpolarized ($\mu_{V}=\mu_{A}=\mu_{H}=0$).
@@
In the limit when the helicity charge is conserved, $\tau_{H}\rightarrow\infty$, the square brackets on the second line of Eq. (58) evaluate to $\cos(c_{h}kt)$ and the integral with respect to $k$ can be performed analytically (see also Eq. (3.36) of Ref. Morales-Tejera:2024uzg ):
$\lim_{\tau_{H}\rightarrow\infty}\delta\bar{\mu}_{V}(t,z)=\frac{\delta\bar{\mu}_{V;0}}{2}$
$\qquad\qquad\left[e^{-(z-c_{h}t)^{2}/2\sigma^{2}}+e^{-(z+c_{h}t)^{2}/2\sigma^{2}}\right],$
$\lim_{\tau_{H}\rightarrow\infty}\delta\bar{\mu}_{H}(t,z)=\frac{\delta\bar{\mu}_{V;0}}{2}$
$\qquad\qquad\left[e^{-(z-c_{h}t)^{2}/2\sigma^{2}}-e^{-(z+c_{h}t)^{2}/2\sigma^{2}}\right].$ (59)
In Fig. 1, we represented $\delta\bar{\mu}_{V}(t,z)$ and $\delta\bar{\mu}_{H}(t,z)$ with respect to $z/\sigma$, when $c_{h}t/\sigma=5$, for the case of a neutral background ($\mu_{V}=\mu_{A}=\mu_{H}=0$). We considered various values of $c_{h}\tau_{H}/\sigma$. It can be seen that, compared to the case when the helicity charge is conserved and $\tau_{H}\rightarrow\infty$ (shown with the black line), the propagation features are damped as $\tau_{H}$ is decreased. At the lowest value of $c_{h}\tau_{H}/\sigma=0.5$, $\delta\mu_{V}(t,z)$ simply decays on the time scale given by $\tau_{H}$.